Using identities, evaluate. 
 297 x 303 ????????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

297 x 303
= (300 - 3 )(300 + 3)
= 300^2- 3^2
= 90000 - 9
= 89991

Thanks